OpenStudy (anonymous):

What is the slope of the line 5x – 2y = 10? -5/2 5/2 2/5 -2/5

OpenStudy (anonymous):

Slope will be : 5/2..

OpenStudy (anonymous):

5/2

OpenStudy (anonymous):

Slope is given by: negative of x coefficient divide by y coefficient..

OpenStudy (diyadiya):

Write it in the form y=mx + b where 'm' is the slope 5x – 2y = 10 -2y= 10 -5x -2y= -5x + 10 y= (-5x)/-2 + 10/-2 y=5/2 x -5 this is of the form y=mx + b so the slope is 5/2

OpenStudy (anonymous):

@Diyadiya don't you think it becomes so lengthy??

OpenStudy (diyadiya):

@waterineyes Well, I know your answer is simpler , i just explained how it is done :)

OpenStudy (anonymous):

\[Slope = \frac{- (x \; Coefficient)}{y Coefficient} = \frac{-5}{-2} = \frac{5}{2}\]

OpenStudy (anonymous):

Yes you can do by comparing with y = mx + c, when a tough equation is given like this one then you can use other method too..

OpenStudy (diyadiya):

:)